OpenStudy (anonymous):

if f(x)=sin^2(x) find f^1(pi/4)

OpenStudy (anonymous):

i think your question is \[f \left( x \right)=\sin ^2x,find ~f \prime \left( \frac{ \pi }{ 4} \right)\]

OpenStudy (anonymous):

find f'(x) then put x=pi/4

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

i did that but im not getting the right answer

OpenStudy (anonymous):

let us together we do it. first find f'(x)

OpenStudy (anonymous):

-9sin(3x)

OpenStudy (anonymous):

f'(x)=2 sin xcos x=sin 2x

OpenStudy (anonymous):

2x=2*pi/4=pi/2 sin pi/2=1

OpenStudy (anonymous):

is the final answer 1?